# Indoor Location Fingerprinting Privacy: A Comprehensive Survey

### v1

@@ 2.1.2. Signal Propagation Time
The distance between the target and the anchor node is also calculated based on the signal propagation time. These methods are typically more accurate than the RSS method. Time of Arrival (ToA) and Time Difference of Arrival (TDoA) are the two well-known methods of distance calculation. ToA-based algorithms require synchronizing between nodes, and TDoA methods were proposed to tackle it. Due to the fact that TDoA takes into account the synchronization of the transmitters, it can somewhat address the problem of synchronization inaccuracy (Do and Yoo, 2014). However, the performance of the ToA/TDoA-based systems is considerably hampered by the NLOS propagation of the signal (Guvenc et al., 2007).
2.1.3. Phase and Angle of Arrival
@@
The Angle of Arrival (AoA) and Phase of Arrival (PoA) approach estimate position by utilizing the angle and the phase that a signal makes with an antenna array (Wielandt et al., 2018). This improved ranging method is used to measure both the angle and the distance. The need for antenna arrays, which makes this approach complex and expensive, is one of its limitations (Sakpere et al., 2017). The signal’s timing of arrival at certain antenna elements may also be used in this approach, although this requires much more complicated hardware and precise calibration.
2.1.4. Channel State Information
@@
Channel State Information (CSI) is another advanced method for calculating distances in indoor positioning. Unlike RSS, which provides only the amplitude value of a received signal, CSI offers a more precise estimation of the received signal’s characteristics across the entire signal bandwidth. This improvement in accuracy is achieved by capturing the channel’s frequency response received by each antenna, making CSI a valuable tool for distance estimation. While CSI often requires multiple antennas to gather this information effectively, it can be employed in a variety of localization systems, including both range-based and range-free approaches (Yang et al., 2013). One notable advantage of CSI is its ability to provide not only signal amplitude but also phase information, enhancing the richness of data available for positioning calculations.”
2.2. Localization Methods
@@ 2.2.1. Multilateration and Trilateration
With the aid of three or more known nodes and the corresponding associated distances, it is a method for determining the position of the unknown node (Carotenuto et al., 2019). Only three well-known nodes are employed in trilateration, which is a specific instance of multilateration. The target node’s location in a two-dimensional space is determined by the intersection of three fictitious circles. However, because of the NLOS effect, which results in significant positioning errors, these circles do not converge at a single location in the real-world indoor environment.
2.2.2. Triangulation
@@
When the angle of arrival is known, it can be applied for positioning precision. It requires at least two anchor nodes and is moderately precise while being less complex (Pomárico-Franquiz et al., 2014). The accuracy of the AoA estimation is crucial to the technique’s location accuracy. The performance of localization can be improved by increasing the number of anchor nodes.
2.2.3. Fingerprinting
@@
Fingerprinting is a widely used indoor location technique that utilizes Wi-Fi, BLE, FM, AM, and ZigBee wireless access technologies (Yan et al., 2018; Uradzinski et al., 2017; Moghtadaiee and Dempster, 2014; Rahman et al., 2017; Faragher and Harle, 2015). As shown in Fig. 3, two phases are involved in the fingerprint-based localization method: offline and online. A radio map is created using the measured data for each recorded position during the training phase using RSS or CSI data acquired at access points (APs) for various known indoor points called reference points (RPs). The Received Signal Strength (RSS) values from all existing APs at known Reference Points (RPs), along with the corresponding $(x,y)$ coordinates of RPs, are stored in the radiomap within LSP. The recorded radiomap is characterized by a dimension of $M\times N$, where $M$ and $N$ represent the number of RPs and APs, respectively. The fingerprint at a given location $(x_{i},y_{i})$ is denoted as ${rss}_{i}^{RP}=[{rss}_{i1},{rss}_{i2},\ldots,{rss}_{iN}]$, where ${rss}_{ij}$ represents the RSS of ${AP}_{j}$ at the ith RP. Moving to the online phase, the user’s location fingerprint is obtained as a vector of RSS values, $S_{u}=[{RSS}_{u1},{RSS}_{u2},\ldots,{RSS}_{uN}]$, where ${RSS}_{uj}$ signifies the RSS of $AP_{j}$ received by the user. This vector is then transmitted to LSP for localization purposes. During this phase, the user’s RSS vector is compared with the previously stored RSS values in the radiomap, and using machine learning methods, LSP estimates the user’s position, $(x,y)$. If more accurately obtained offline data is used to build the radio map, this method offers great accuracy. However, it takes a great deal of work to create the radio map for large-area deployment (e.g., manpower, time, and cost). Additionally, for dynamic networks, the offline database needs to be regenerated whenever a node’s position, even a single node, is unexpectedly changed or removed.
Figure 3. The flow of indoor location fingerprinting technique.
@@
ML algorithms are frequently used to improve fingerprinting accuracy and aid in building radio maps (Nessa et al., 2020). The simplest algorithm used for fingerprint-based localization techniques is k-Nearest Neighbors (kNN). This algorithm computes a distance metric that measures the separations between the target’s measurements at various APs and the measurements taken during the training phase. Euclidean distance is the most widely used measurement of distance. The radio map’s k closest RPs to the target with the smallest distances are chosen in this algorithm. The target’s position is then inferred by averaging these RPs’ coordinates. However, the kNN algorithm also uses alternative distance measures such as the Manhattan, Mahalanobis, and Minkowski distances (Farshad et al., 2013).
2.2.4. Range-free
@@
Instead of calculating the position based on distance, range-free localization methods are proposed:
@@
  • Centroid: Instead of using distance or angle measurement, this method estimates the location of the unknown node using a geometric relation. Once a reliable communication link has been established between each anchor node and the unknown target node, the positions of the anchor nodes are known. The centroid of the geometric shape formed by the position of the anchor nodes connected to the target node is taken to be the location of the unknown nodes. In (Chen et al., 2013), a Weighted Centroid Localization (WCL) approach-based BLE beacon IPS was suggested.
@@
  • Distance Vector Hop (DV Hop): In a multi-hop setting, this method involves estimating the distance vector depending on the hop count. An information table is kept up to date with the coordinates of the ith node and the minimum hop count value from the anchor node to the ith node. The anchor node broadcasts the position data to the nearby nodes, which subsequently broadcast it to more nodes, and so on. Finding the hop size for a specific hop is a crucial issue for this strategy. The distance between the node that is m hops distant from the anchor node is easily estimated after obtaining the average hop size (Chen et al., 2013).
2.3. Technologies
@@ 2.3.1. Radio Frequency
Radio frequency (RF) technologies are frequently deployed for indoor localization due to their versatility and robust performance in various settings. Among the widely utilized RF technologies are Wi-Fi-based localization systems, which capitalize on the ubiquity of Wi-Fi access points in indoor spaces. These systems employ signal strength and triangulation techniques to estimate the position of Wi-Fi-enabled devices such as smartphones or tablets (Yang and Shao, 2015). Bluetooth Low Energy (BLE) beacons also play a crucial role, especially in asset tracking and proximity-based applications. BLE beacons emit low-power signals that can be detected by smartphones or dedicated receivers, facilitating accurate indoor positioning (Faragher and Harle, 2014). Ultra-wideband (UWB) technology, recognized for its high precision, is gaining traction in applications demanding centimeter-level accuracy, such as industrial robotics and healthcare (Alarifi et al., 2016). FM/AM infrastructures, from another perspective, can be used to deploy various indoor localization methods without the need for additional hardware, which is applicable for long-range solutions with susceptibility to interference  (Moghtadaiee and Dempster, 2014; Rahman et al., 2017).
2.3.2. Non-Radio Frequency
@@
Conversely, non-radio frequency wireless technologies are emerging as alternatives to RF solutions. Infrared (IR) technology relies on the transmission of light signals in the infrared spectrum to determine the location of objects equipped with IR transmitters and receivers. Although less common than RF technologies, IR systems offer advantages in specific scenarios, such as indoor positioning within controlled environments or line-of-sight requirements. Additionally, visible light communication (VLC) leverages LED lighting infrastructure to transmit data and positional information through modulated light signals. VLC systems have demonstrated potential for indoor localization while providing energy-efficient illumination (Alam et al., 2020).
3. Indoor Localization Applications and Privacy Concerns
3.1. Marketing
Contextually-aware location-based marketing strategies have the potential to significantly boost sales and profits in e-commerce by delivering tailored marketing directly to consumers based on their location, especially in shopping centers, utilizing location-based technology. This approach allows real-time communication between consumers and sellers, enhancing the overall shopping experience. Additionally, it aids businesses in tracking customer behaviors, patterns, and foot traffic, offering personalized services based on mobile devices’ information (Nessa et al., 2020; Singh et al., 2021). However, these strategies bring privacy challenges, as the use of location-based technology for personalized marketing raises concerns about user privacy and potential misuse of real-time location data. Tracking customer behaviors without explicit consent and the extensive collection of personal data through personalized mobile devices also pose ethical and privacy issues. Furthermore, the precision of IPS in localizing mobile devices indoors and outdoors intensifies concerns about intrusive tracking if not appropriately safeguarded.
3.2. Ambient Assisted Living and Disaster Management
Ambient assisted living platforms, relying on precise indoor location tracking through technologies like Bluetooth and various IPS, offer significant benefits for the elderly, ill, or disabled individuals, particularly those with neurodegenerative diseases (Wan et al., 2015). These platforms facilitate behavioral tracking, monitoring daily activities, movement patterns, vital signs, and detecting potential dangers such as falls or injuries. Smart houses leverage indoor localization to enhance user experiences, enabling the homeowner to control Wi-Fi network access based on device presence (Bonafini et al., 2019). Beyond homes, location tracking techniques find applications in agriculture for monitoring greenhouses, libraries for book location, parking garages for car monitoring, and warehouses for item tracking (Singh et al., 2021). In disaster management, indoor localization can pinpoint the precise location of individuals in peril for efficient rescue efforts (Chehri et al., 2011; Asimakopoulou and Bessis, 2011). Despite these advantages, privacy concerns arise, including the continuous monitoring of sensitive health-related data, potential invasions of privacy in smart homes, perceived intrusive surveillance in various contexts, and privacy implications in disaster management. Striking a balance between the benefits and privacy protection is crucial, necessitating robust data safeguards and ethical guidelines.
3.3. Health Services and Public Safety
Indoor localization holds immense potential for elevating service standards in healthcare, aiding hospital staff in swiftly locating patients and enabling patients to navigate therapy rooms independently. Doctors can monitor patient mobility, visitors can easily find their patients, and operating rooms can be well-equipped (Singh et al., 2021; Calderoni et al., 2015). Smart devices’ positions are traceable, facilitating their quick retrieval. The technology’s future applications include nanosensor-based drug delivery for tumor targeting and hazard identification for disaster mitigation (Giuliano et al., 2013). In law enforcement, Bluetooth beacon-based systems aid quick responses, while in smart buildings, Wi-Fi-enabled alarms guide people to safety. Despite these benefits, concerns arise about continuous tracking compromising patient privacy in healthcare, nanosensor-based drug delivery raising intimate monitoring questions, and surveillance implications in safety management and law enforcement. Striking a balance requires careful ethical and legal considerations in these applications.
During the COVID-19 pandemic, for example, IPS were employed to monitor and enforce social distancing guidelines in various settings, including hospitals and workplaces. While crucial for public health, these technologies raise significant privacy concerns. The continuous and precise tracking of individuals’ movements within indoor spaces may result in the collection of sensitive location data, revealing interactions, routines, and patterns. This raises potential privacy infringements, and there’s a risk of unauthorized access or misuse, challenging data security and individual autonomy. Addressing this complex challenge requires robust privacy safeguards, transparent data handling practices, and clear regulations to ensure the responsible deployment of IPS for COVID-19 measures while respecting privacy rights and maintaining public trust.
3.4. Surveillance and Tracking
LBS extend beyond human tracking, finding significant applications in industrial settings, where the automatic tracking of numerous objects is essential for effective management. This involves real-time control of item locations alongside localization and identification, requiring new Medium Access Control layer protocols to prevent collisions (Li, 2021). The prediction of autonomous robot locations, especially in the emerging era of collaborative robots (co-bots), enhances safety and efficiency in various tasks, such as safety and collision avoidance in industrial duties (Bentley et al., 2014; Chekuri and Won, 2017; Sadowski and Spachos, 2018). However, the continuous monitoring of objects and devices in industrial contexts raises privacy concerns for workers, impacting workplace privacy. Anticipating robot locations and monitoring employee interactions could similarly affect privacy. Furthermore, the use of location-based access control and continuous tracking for anti-theft systems raises data security and surveillance issues. Additionally, indoor localization in entertainment and broadcasting may involve tracking event attendees, presenting privacy challenges. Achieving a balance between the advantages of these technologies and privacy protection requires the establishment of robust safeguards and guidelines for responsible use.
4. Privacy Leakage In Indoor Location Fingerprinting
@@
This section provides an in-depth examination of the indoor location fingerprinting system emphasizing privacy vulnerabilities and leakages. We provide a comprehensive understanding of the system entities and the underlying data structure, particularly in the context of employing location fingerprinting techniques for indoor localization. Through this, we derive insights into potential areas of privacy leakage within these systems. Additionally, we extend this section to a detailed exploration of the broader privacy leakage arising from the information inferred about the users in indoor location fingerprinting.
4.1. Entities-Driven Privacy Leakage
@@
Here, we introduce and analyze the fundamental components associated with the processes of data transmission and service provisioning in the indoor location fingerprinting system so that we can comprehend the potential privacy vulnerabilities. As shown in Fig. 4, the entities include Access Points (APs), Users, Trusted Third Parties (TTPs), LBS servers, Content Providers, and Data Publishers. For each entity, we first introduce their characteristics, followed by the inherent privacy leakages within each entity, and then the potential privacy vulnerabilities that these entities may introduce.
Figure 4. Privacy-vulnerable entities in indoor location fingerprinting systems.
4.1.1. Access Point (AP)
@@
The potential privacy leakage for APs are the disclosure of their locations. The transmission of RSS/CSI data by APs to users during the service stage can potentially expose APs’ locations through various methods. The disclosure of this data for APs, particularly through RSS, makes them susceptible to malicious activities, leading to physical access and potential damage. The distance from an AP can be calculated using RSS information, and multilateration and triangulation techniques can be applied to pinpoint the AP’s location. Researchers have also found that the ratio of signal powers between the first peak and second peak in CSI can be used as a physical layer metric to estimate the distance from the AP (Wang et al., 2015).
APs can be also curious and pose a privacy threat themselves. They are indispensable in IPS, but they have less power than a malicious localization server. The user collects one measuring value from each AP during localization. Theoretically, localization requires at least three measuring values for a 2-dimensional area, (or four for a 3D area) which are located in different places. If the communication during measurement is unilateral, the measuring result is inherently secure against curious APs. Take RSS as an example, RSS values are measured at the user device by evaluating the signal strength from each AP without sending any information back. This kind of measurement leaks no information to the APs. If the communication during measurement is bilateral, colluding APs can obtain the full measuring result of the user and compute the location with the help of the localization server. Hence, it is crucial to make the measurements unlinkable or anonymous for measuring techniques involving bilateral communications.
4.1.2. Users

### v2

@@ 2.1.2 Signal Propagation Time
The distance between the target and the anchor node is also calculated based on the signal propagation time. These methods are typically more accurate than the RSS method. Time of Arrival (ToA) and Time Difference of Arrival (TDoA) are well-known distance calculation methods. ToA-based algorithms require synchronizing between nodes, and TDoA methods were proposed to tackle it. Since TDoA takes into account the synchronization of the transmitters, it can somewhat address the problem of synchronization inaccuracy [19]. However, the performance of the ToA/TDoA-based systems is considerably hampered by the NLOS propagation of the signal [20].
2.1.3 Phase and Angle of Arrival
@@
The Angle of Arrival (AoA) and Phase of Arrival (PoA) approach estimate position by utilizing the angle and the phase that a signal makes with an antenna array [21]. This improved ranging method is used to measure both the angle and the distance. The need for antenna arrays, which makes this approach complex and expensive, is one of its limitations [22]. The signal’s timing of arrival at certain antenna elements may also be used in this approach, although this requires much more complicated hardware and precise calibration.
2.1.4 Channel State Information
@@
Channel State Information (CSI) is another advanced method for calculating distances in indoor positioning. Unlike RSS, which provides only the amplitude value of a received signal, CSI offers a more precise estimation of the received signal’s characteristics across the entire signal bandwidth. This improvement in accuracy is achieved by capturing the channel’s frequency response received by each antenna, making CSI a valuable tool for distance estimation. While CSI often requires multiple antennas to gather this information effectively, it can be employed in a variety of localization systems, including both range-based and range-free approaches [23]. One notable advantage of CSI is its ability to provide not only signal amplitude but also phase information, enhancing the richness of data available for positioning calculations.”
2.2 Localization Methods
@@ 2.2.1 Multilateration and Trilateration
With the aid of three or more known nodes and the corresponding associated distances, it is a method for determining the position of the unknown node [24]. Only three well-known nodes are employed in trilateration, which is a specific instance of multilateration. The target node’s location in a two-dimensional space is determined by the intersection of three fictitious circles. However, because of the NLOS effect, which results in significant positioning errors, these circles do not converge at a single location in the real-world indoor environment.
2.2.2 Triangulation
@@
When the angle of arrival is known, it can be applied for positioning precision. It requires at least two anchor nodes and is moderately precise while being less complex [25]. The accuracy of the AoA estimation is crucial to the technique’s location accuracy. The performance of localization can be improved by increasing the number of anchor nodes.
2.2.3 Fingerprinting
@@
Fingerprinting is a widely used indoor location technique that utilizes Wi-Fi, BLE, FM, AM, and ZigBee wireless access technologies [26, 27, 4, 5, 28]. As shown in Fig. 3, two phases are involved in the fingerprint-based localization method: offline and online. A radio map is created using the measured data for each recorded position during the offline phase using RSS or CSI data acquired at access points (APs) for various known indoor points called reference points (RPs). The Received Signal Strength (RSS) values from all existing APs at known Reference Points (RPs), along with the corresponding $(x,y)$ coordinates of RPs, are stored in the radiomap within LSP. The recorded radiomap is characterized by a dimension of $M\times N$, where $M$ and $N$ represent the number of RPs and APs, respectively. The fingerprint at a given location $(x_{i},y_{i})$ is denoted as ${rss}_{i}^{RP}=[{rss}_{i1},{rss}_{i2},\ldots,{rss}_{iN}]$, where ${rss}_{ij}$ represents the RSS of ${AP}_{j}$ at the ith RP. Moving to the online phase, the user’s location fingerprint is obtained as a vector of RSS values, $S_{u}=[{RSS}_{u1},{RSS}_{u2},\ldots,{RSS}_{uN}]$, where ${RSS}_{uj}$ signifies the RSS of $AP_{j}$ received by the user. This vector is then transmitted to LSP for localization purposes. During this phase, the user’s RSS vector is compared with the previously stored RSS values in the radiomap, and using machine learning methods, LSP estimates the user’s position, $(x,y)$. If more accurately obtained offline data is used to build the radio map, this method offers great accuracy. However, it takes a great deal of work to create the radio map for large-area deployment (e.g., manpower, time, and cost). Additionally, for dynamic networks, the offline database needs to be regenerated whenever a node’s position, even a single node, is unexpectedly changed or removed.
Figure 3: The flow of ILF technique.
@@
ML algorithms are frequently used to improve fingerprinting accuracy and aid in building radio maps [9]. The simplest algorithm used for fingerprint-based localization techniques is k-Nearest Neighbors (kNN). This algorithm computes a distance metric that measures the separations between the target’s measurements at various APs and the measurements taken during the training phase. Euclidean distance is the most widely used measurement of distance. The radio map’s k closest RPs to the target with the smallest distances are chosen in this algorithm. The target’s position is then inferred by averaging these RPs’ coordinates. However, the kNN algorithm can also use alternative distance measures such as the Manhattan, Mahalanobis, and Minkowski distances [29].
2.2.4 Range-free
@@
Instead of calculating the position based on distance, range-free localization methods were proposed:
@@
  • Centroid: Instead of using distance or angle measurement, this method estimates the location of the unknown node using a geometric relation. Once a reliable communication link has been established between each anchor node and the unknown target node, the positions of the anchor nodes are known. The centroid of the geometric shape formed by the position of the anchor nodes connected to the target node is taken to be the location of the unknown nodes. In [30], a Weighted Centroid Localization (WCL) approach-based BLE beacon IPS was suggested.
@@
  • Distance Vector Hop (DV Hop): In a multi-hop setting, this method involves estimating the distance vector depending on the hop count. An information table is kept up to date with the coordinates of the ith node and the minimum hop count value from the anchor node to the ith node. The anchor node broadcasts the position data to the nearby nodes, which subsequently broadcast it to more nodes, and so on. Finding the hop size for a specific hop is a crucial issue for this strategy. The distance between the node that is m hops distant from the anchor node is easily estimated after obtaining the average hop size [30].
2.3 Technologies
@@ 2.3.1 Radio Frequency
Radio frequency (RF) technologies are frequently deployed for indoor localization due to their versatility and robust performance in various settings. Among the widely utilized RF technologies are Wi-Fi-based localization systems, which capitalize on the ubiquity of Wi-Fi access points in indoor spaces. These systems employ signal strength and triangulation techniques to estimate the position of Wi-Fi-enabled devices such as smartphones or tablets [31]. Bluetooth Low Energy (BLE) beacons also play a crucial role, especially in asset tracking and proximity-based applications. BLE beacons emit low-power signals that can be detected by smartphones or dedicated receivers, facilitating accurate indoor positioning [32]. Ultra-wideband (UWB) technology, recognized for its high precision, is gaining traction in applications demanding centimeter-level accuracy, such as industrial robotics and healthcare [33]. FM/AM infrastructures, from another perspective, can be used to deploy various indoor localization methods without the need for additional hardware, which is applicable for long-range solutions with susceptibility to interference  [4, 5].
2.3.2 Non-Radio Frequency
@@
Conversely, non-radio frequency wireless technologies are emerging as alternatives to RF solutions. Infrared (IR) technology relies on the transmission of light signals in the infrared spectrum to determine the location of objects equipped with IR transmitters and receivers. Although less common than RF technologies, IR systems offer advantages in specific scenarios, such as indoor positioning within controlled environments or line-of-sight requirements. Additionally, visible light communication (VLC) leverages LED lighting infrastructure to transmit data and positional information through modulated light signals. VLC systems have demonstrated potential for indoor localization while providing energy-efficient illumination [34].
3 Privacy Leakage In Indoor Location Fingerprinting
@@
This section provides an in-depth examination of the ILF system emphasizing privacy vulnerabilities and leakages. We offer a comprehensive understanding of the system entities and the underlying data structure, particularly in employing location fingerprinting techniques for indoor localization. Through this, we derive insights into potential areas of privacy leakage within these systems. Additionally, we extend this section to a detailed exploration of the broader privacy leakage arising from the information inferred about the users in ILF.
3.1 Entities-Driven Privacy Leakage
@@
Here, we introduce and analyze the fundamental components associated with the processes of data transmission and service provisioning in the ILF system so that we can comprehend the potential privacy vulnerabilities. As shown in Fig. 4, the entities include Access Points (APs), Users, Trusted Third Parties (TTPs), Location Service Providers (LSPs), Content Providers, and Data Publishers. For each entity, we first introduce their characteristics, followed by the inherent privacy leakages within each entity, and then the potential privacy vulnerabilities that these entities may introduce.
Figure 4: Privacy-vulnerable entities in ILF systems.
3.1.1 Access Point (AP)
@@
The potential privacy leakage for APs is the disclosure of their locations. The transmission of RSS/CSI data by APs to users during the service stage can potentially expose APs’ locations through various methods. The disclosure of this data for APs, particularly through RSS, makes them susceptible to malicious activities, leading to physical access and potential damage. The distance from an AP can be calculated using RSS information, and multilateration and triangulation techniques can be applied to pinpoint the AP’s location. Researchers have also found that the ratio of signal powers between the first peak and second peak in CSI can be used as a physical layer metric to estimate the distance from the AP [35].
APs can be also curious and pose a privacy threat themselves. They are indispensable in IPS, but they have less power than a malicious localization server. The user collects one measuring value from each AP during localization. Theoretically, localization requires at least three measuring values for a 2-dimensional area, (or four for a 3D area) which are located in different places. If the communication during measurement is unilateral, the measuring result is inherently secure against curious APs. Take RSS as an example, RSS values are measured at the user device by evaluating the signal strength from each AP without sending any information back. This kind of measurement leaks no information to the APs. If the communication during measurement is bilateral, colluding APs can obtain the full measuring result of the user and compute the location with the help of the localization server. Hence, making the measurements unlinkable or anonymous for measuring techniques involving bilateral communications is crucial. From another perspective, when transmitting data, the user’s device will automatically connect to the nearest AP for communication. As a result, this AP becomes aware of the user’s proximity, which can inadvertently reveal the user’s location to the AP, potentially compromising privacy. This connection process implies that even without detailed data exchange, the mere act of connecting provides the AP with information about the user’s presence in its vicinity.
3.1.2 Users